OpenStudy (anonymous):

Write the following summation as a definite integral and then evaluate using the Fundamental Theorem. Your answer must show the definite integral, the antiderivative, and the exact answer. Where do I even begin?

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} \frac{ 3 }{ n } \sum_{i=1}^{n} \left( 1+\left( \frac{ 3i }{ n } \right)^{2} \right)\]